OpenStudy (anonymous):

(1/x+h) + (1/x) -------------- x

OpenStudy (anonymous):

\[\frac{1}{x(x+h)}+\frac{1}{x^2}\]

hartnn (hartnn):

which if u want can also be written as \[\frac{2x+h}{(x^2)(x+h)}\]

OpenStudy (anonymous):

is it possible for you to show me that step by step?

hartnn (hartnn):